OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 3x − 7 3x − y = −2

Directrix (directrix):

y = x2 + 3x − 7 3x − y = −2 Solve for y in the second equation. y = 3x + 2 In the first equation, substitute 3x + 2 for y. 3x + 2 = x2 + 3x − 7 and solve for x. Then, you'll need to find the y-values that go with each of those x-values.

Directrix (directrix):

@mabitrix ^^^